OpenStudy (anonymous):

Pure acid is to be added to a 10% acid solution to obtain 99L of a 20% acid solution. What amounts of each should be used? (answer 1: L of pure acid solution, answer 2: L of 10% acid solution)

jimthompson5910 (jim_thompson5910):

Let x = amount of 100% acid solution (pure acid, no water) y = amount of 10% acid solution (acid + water) both amounts are in liters

jimthompson5910 (jim_thompson5910):

we have a solution that is 10% acid. We have y liters of this solution, so we have 0.10y liters of pure acid with this 10% solution (the rest is water) x liters of pure acid are added to this to get x + 0.10y liters total of pure acid

jimthompson5910 (jim_thompson5910):

99L of 20% acid ---> 0.20*99 = 19.8 so 19.8 liters of the final solution are pure acid

jimthompson5910 (jim_thompson5910):

this means x + 0.10y = 19.8 is one equation

jimthompson5910 (jim_thompson5910):

the two amounts x & y must add to 99 because this is the total amount of liters in the final solution of water and acid x+y = 99

jimthompson5910 (jim_thompson5910):

solve x+y = 99 for y to get y = 99-x then plug that into the first equation x + 0.10y = 19.8 x + 0.10(99-x) = 19.8 I'll let you finish up